Motorcycled travelled 90kph approaches a car traveling in the same direction at 75kph. When the motor cycle is 60m behind the car the rider pushes down on the acceleration and passes by the car 10sec. later. what is the acceleration of a motor cycle?
1
Expert's answer
2019-01-14T08:59:11-0500

In case of uniformly accelerated motion the distance could be found by the following formula:

S = V0t + 〖at〗^2/2, where V0 is the velocity difference of the car and the motorcycle or the approach speed.

Find a:

a = (2(S-V_0 t))/t^2 , where V0 = 15 kph or 4.2 m/s

a = (2×(60-4.2×10))/100 = 0.36 (m/s2)